solve the system of x-2y=2 x+y=5
Answer by hamsanash1981@gmail.com(151)

x - 2y = 2 ---1
x + y = 5 ---2
multiplying equation 2 by 2
2x+2y = 10 ---3
solving equ 1 and 3
x - 2y = 2
2x + 2y = 10
-----------
3x = 12
x= 4
then y = 5-4
y= 1
therefore x=4 and y = 1